# Rambo: Enhancing RAG-based Repository-Level Method Body Completion

### v1

@@ 3 Mono<Void> save(Mono<RouteDefinition> route) {
Fig. 3 shows Rambo’s performance across different reference sizes (LOC). As expected, Rambo performs much better with smaller code snippets (1–50 LOC), achieving high CodeBLEU, CR, and EM. However, Rambo’s performance declines gracefully as the reference size increases, particularly for larger codebases (201+ LOC). This trend also appears in the results of the other approaches [26]. This decline can be attributed to the increased complexity and dependencies within larger code, which challenge Rambo’s ability to maintain syntactical and functional accuracy. These findings suggest that while Rambo is highly effective for smaller tasks, improvements are needed to enhance its scalability and performance on larger, more complex codebases.
@@
In Table II, the experimental results show that the repo-level MBC approaches perform much better on Defect4J than Rambo’s benchmark. This discrepancy is attributed to the fact that Rambo’s benchmark includes larger repositories and more complex MBC problems with longer reference codes. Additionally, the MBC problems in Rambo’s benchmark demand an average of 7.2 repo-specific code element invocations (with a maximum of 43), whereas the corresponding problems in Defects4J require only about 4.8 invocations (with a maximum of 17).
@@ 5.2.1 Impact of Retrieval Technique
TABLE III: Performance of Rambo by Retrieval Techniques
@@
Table III compares the performance of Rambo across three different retrieval methods. As seen, the performance of Rambo is slightly impacted by the retrieval techniques with the relative margins of 1.6% in CodeBLEU, 2.4% in CR, and 4.8% in EM. Although Semantic Retrieval offers broader contextual insights, its performance slightly lags behind the other techniques. The reason could be that Rambo primarily relies on identifying functionally similar methods and code elements by comparing method signatures and element names, where deep semantic understanding may not be as critical. The Hybrid approach, which combines both lexical and semantic methods, marginally outperforms the individual approaches but introduces additional complexity. Given the small performance gains, this added complexity may not be justified compared to the efficiency of lexical retrieval. Indeed, Lexical retrieval, which focuses on token-level similarity, stands out for its balance of performance and efficiency. Its lightweight design allows for fast and accurate retrieval of relevant code snippets, making it particularly effective for Rambo’s MBC tasks.
@@ 5.2.2 EEI and RUE: Contribution Analysis
To evaluate the impact of Essential Code Element Identification (EEI) and Relevant Usage Extraction (RUE) on Rambo’s overall performance, we developed two variants of Rambo. The first variant contains only the component of Sketch generation serving as a baseline for evaluating performance without the additional contextual information provided by EEI or RUE. The second variant, Essential (code) elements only, analyzes the sketch code to identify most similar elements in the repository and uses their signatures as well as their bodies as the context for code generation by $\mathscr{M}$.
Table IV shows the performance of these two variants compared to Rambo. The Sketch generation component serves as a baseline, achieving a CodeBLEU of 49.75, a CR of 61.06, and an EM score of 14.04. Interestingly, the performance of the Essential elements only variant is lower than the sketch-only approach. Although the essential elements and their bodies are included as context, the performance decrease suggests that introducing them without proper usage context (as in RUE) may mislead the model. The retrieval of functionally similar code elements alone is insufficient without understanding their specific role in the repository, leading to less accurate code generation.
@@
TABLE IV: Contributions of EEI and RUE on Rambo’s performance
@@
Essential (code) elements only 41.10 61.49 13.40
@@ 5.2.3 Impact of Relevant Context’s Size
Fig. 4 shows the performance of Rambo as the size of the relevant context, $C_{R}$, increases, measured by the number of relevant methods included. The performance of Rambo improves consistently as more relevant context is incorporated, with notable gains of 59% in CodeBLEU, 37% in Compilation Rate (CR), and a 5X increase in Exact Match (EM) when the context size increases from 0 to 20 relevant methods. The most substantial improvement occurs when the context size grows from 0 to 5 methods. However, after a certain threshold, including additional relevant methods in $C_{R}$ results in diminishing returns. This suggests that Rambo can effectively extract essential information with a relatively small set of relevant methods, and beyond that, further context offers marginal benefits.
@@ 5.2.4 Impact of Code LLM’s Size
Figure 5: Rambo’s performance by the sizes of $\mathscr{M}_{g}$ (DeepSeek Coder)
{mdframed}
[backgroundcolor=light-gray, skipabove=5pt, skipbelow=0pt, nobreak=false ]{mdtheorem}namelabel
Answer to RQ2: Rambo’s performance improves with relevant context and larger Code LLMs. Lexical retrieval provides the best trade-off between performance and efficiency. Performance gains diminish with excessive relevant context ($C_{R}$) and larger models increase accuracy but slow down generation.
@@ 5.3.2 Impact of Repository Size
Fig. 7 illustrates Rambo’s performance across repositories of various sizes. Rambo performs optimally on smaller repositories (20 out of 26 repositories with fewer than 800 files). On average, it achieves a CodeBLEU score of 54.5, a Compilation Rate (CR) of 64%, and an Exact Match (EM) rate of 21.1%, with generation times remaining under 2 seconds. However, as the repository size increases, there is a decline in performance across all metrics. In particular, for very large repositories containing more than 2,000 files, CodeBLEU drops to 49.23, EM decreases to 13.95%, and generation times increase significantly. These findings suggest that while Rambo excels in smaller repositories, its ability to maintain high accuracy and efficiency diminishes with larger, more complex repositories, where longer generation times and lower performance are observed.
Figure 7: Rambo’s performance by repository sizes (files), left axis: CodeBLEU, CR, and EM; right axis: Code gen. time (s)
{mdframed}
[backgroundcolor=light-gray, skipabove=5pt, skipbelow=0pt, nobreak=false ]{mdtheorem}namelabel Answer to RQ3: Rambo performs best with balanced surrounding context ($C_{f}$) and excels in smaller repositories. Performance declines in larger repositories due to increased complexity and generation times.
@@ 5.5 Threats to Validility
Threats to internal validity include the hyperparameter settings of both the baseline models and Rambo. To mitigate this threat, we systematically varied the settings of Rambo and thoroughly examined its performance (as discussed in Section 5). Additionally, we reused the implementations and settings from the original papers of the baseline approaches [15] to ensure consistency in comparison. Another potential threat arises from the methods used to identify essential code elements within the repository. To address this, we employed a robust retrieval mechanism informed by existing studies, ensuring our approach aligned with established practices. Furthermore, to minimize the threat of implementation errors, we conducted a detailed review of our code and have made it publicly available [26] to facilitate replication and validation by other researchers.
Threats to construct validity relate to the appropriateness of our evaluation procedure. We utilized widely accepted metrics for both text-based and code-based evaluations, including CodeBLEU, Exact Match, Compilation Rate, and pass@k, which are standard measures in code completion and code generation research. A potential threat remains due to the absence of a human evaluation, which could provide a more nuanced assessment of our approach. To address this limitation, we plan to incorporate human evaluations in future work, wherein experienced developers will assess the generated method bodies based on criteria such as functionality, accuracy, and adherence to repository standards. Additionally, the controlled environments used in our experiments may not fully represent real-world scenarios. To mitigate this, we conducted evaluations in diverse settings and examined performance across multiple scenarios and configurations.
Threats to external validity primarily concern the generalizability of our results to different repositories and programming languages. To reduce this threat, we selected representative code LLMs models and repositories widely recognized in both NLP and SE communities [7, 19, 6]. Another potential threat is using a limited set of Java repositories in our experiments. Our current study focuses on a specific set of repositories, which may limit the applicability of our findings to other types of codebases or programming languages. In future work, we intend to extend our experiments to include a broader range of repositories and languages to enhance the generalizability of our results.
@@ 6 Related Work
Rambo is closely related to the research on repository-level code completion, which utilizes the broader context provided by an entire code repository [32, 33, 34, 35, 15, 16, 18]. RepoPrompts [35] introduces an approach that first generates prompts reflecting the repository’s structure and predefined relevant contexts, then learns to select the prompt most likely to produce the desired output. In the work by Wang et al.  [18], the task of retrieving useful context is framed as a reinforcement learning problem, where the retriever parameters are iteratively updated. MGD [32] monitors the code generated by LLMs, queries static analysis tools in the background, and leverages the resulting information to guide the LLMs during the decoding stage. RepoCoder [15] integrates a similarity-based retriever and a code LLM into an iterative retrieval-generation pipeline. Although these efforts show promising performance, Rambo differs from them in that Rambo does not require learning/training or touching code LLMs. Additionally, instead of similarity-based retrieve, Rambo emphasizes the extraction of code elements most relevant to the infilling method and their usages, ensuring that the generated method body is both contextually accurate and aligned with the repository’s specific coding practices.
@@
Several learning-based approaches have been proposed for specific SE tasks such as code translation/migration [47, 48, 49, 50, 51, 52], text-code translation [53, 54, 55, 56, 57], test generation [58, 59], bug detection [60, 61, 62, 63], bug fix identification [64], and program repair [65, 66].
@@ 7 Conclusion
In this paper, we introduced Rambo, a novel Retrieval-Augmented Generation (RAG)-based approach designed specifically for repository-level method body completion (MBC). Our approach centers on the accurate identification and utilization of essential repo-specific code elements, significantly enhancing the performance of code LLMs in generating contextually relevant and syntactically correct method bodies. Our experimental results show Rambo’s superiority across all key metrics, including CodeBLEU, Compilation Rate, pass@k, and Exact Match. Notably, Rambo not only surpassed existing methods but also exceeded the performance of RepoCoder Oracle, setting a new benchmark in the field. These show Rambo’s robustness, accuracy, and practical utility in real-world code generation tasks. By leveraging essential code elements and relevant usages, Rambo addresses the limitations of existing similarity-based approaches, providing a more precise and reliable solution for method body completion in complex repositories.

### v3

@@ 3 Mono<Void> save(Mono<RouteDefinition> route) {
Fig. 3 shows Rambo’s performance across different reference sizes (LOC). As expected, Rambo performs much better with smaller code snippets (1–50 LOC), achieving high CodeBLEU, CR, and EM. However, Rambo’s performance declines gracefully as the reference size increases, particularly for larger codebases (201+ LOC). This trend also appears in the results of the other approaches [26]. This decline can be attributed to the increased complexity and dependencies within larger code, which challenge Rambo’s ability to maintain its accuracy. These findings suggest that while Rambo is highly effective for smaller tasks, improvements are needed to enhance its scalability and performance on larger, more complex codebases.
@@
In Tables 3 and 2, the experimental results show that the repo-level MBC approaches perform much better on Defect4J than Rambo’s benchmark. This discrepancy is attributed to the fact that Rambo’s benchmark includes larger repositories and more complex MBC problems with longer reference codes. Additionally, the MBC problems in Rambo’s benchmark demand an average of 7.2 repo-specific code element invocations (with a maximum of 43), whereas the corresponding problems in Defects4J require only about 4.8 invocations (with a maximum of 17).
@@ 5.2.1 Impact of Retrieval Technique
Table 4: Performance of Rambo by Retrieval Techniques
@@
Table 4 compares the performance of Rambo across three different retrieval methods. As seen, the performance of Rambo is slightly impacted by the retrieval techniques with the relative margins of 1.6% in CodeBLEU, 2.4% in CR, and 4.8% in EM. Although Semantic Retrieval offers broader contextual insights, its performance slightly lags behind the other techniques. The reason could be that Rambo primarily relies on identifying functionally similar methods and code elements by comparing method signatures and element names, where deep semantic understanding may not be as critical. The Hybrid approach, which combines both lexical and semantic methods, marginally outperforms the individual approaches but introduces additional complexity. Given the small performance gains, this added complexity may not be justified compared to the efficiency of lexical retrieval. Indeed, Lexical retrieval, which focuses on token-level similarity, stands out for its balance of performance and efficiency. Its lightweight design allows for fast and accurate retrieval of relevant code snippets, making it particularly effective for Rambo’s MBC tasks.
@@ 5.2.2 EEI and RUE: Contribution Analysis
To evaluate the impact of Essential Code Element Identification (EEI) and Relevant Usage Extraction (RUE) on Rambo’s overall performance, we built two variants of Rambo. The first one contains only the component of Sketch generation serving as a baseline for evaluating performance without the additional contextual information provided by EEI or RUE. The second variant, Essential (code) elements only, analyzes the sketch code to identify most similar elements in the repository and uses their signatures as well as their bodies as the context for code generation by $\mathscr{M}$.
Table 5 shows the performance of these two variants compared to Rambo. The Sketch generation component serves as a baseline, achieving a CodeBLEU of 49.75, a CR of 61.06, and an EM score of 14.04. Interestingly, the performance of the Essential elements only variant is lower than the sketch-only approach. Although the essential elements and their bodies are included as context, the performance decrease suggests that introducing them without proper usage context (as in RUE) may mislead the model. The retrieval of functionally similar code elements alone is insufficient without understanding their specific role in the repository, leading to less accurate code generation.
@@
Table 5: Contributions of EEI and RUE on Rambo’s performance
@@
Essential elements only 41.10 61.49 13.40
@@
5.2.3 Impact of Number of Sketches
To study the impact of sketch generation on Rambo’s performance, we varied the number of generated sketches in Sketch Generation of Rambo on Defects4J with DeepSeek Coder as the code LLM. When using a single sketch, Rambo achieves a BLEU score of 51.03, an ES of 53.05, and an EM rate of 23.06. Increasing the number of sketches to five results in lower scores, e.g., BLEU of 49.43, CodeBLEU of 51.12, and EM of 21.57 in the case of generating 5 sketches. In our experiments, we found that that the sketches generated are not substantially different from one another. Additionally, the sketches may even introduce irrelevant elements that slightly degrade performance. Consequently, using a single sketch is sufficient for effectively identifying the essential code elements for method body completion.
@@ 5.2.4 Impact of Relevant Context’s Size
Fig. 4 shows the performance of Rambo as the size of the relevant context, $C_{R}$, increases, measured by the number of relevant methods included. The performance of Rambo improves consistently as more relevant context is incorporated, with notable gains of 59% in CodeBLEU, 37% in Compilation Rate, and a 5X increase in Exact Match when the context size increases from 0 to 20 relevant methods. The most substantial improvement occurs when the context size grows from 0 to 5 methods. However, after a certain threshold, including additional relevant methods in $C_{R}$ results in diminishing returns. This suggests that Rambo can effectively extract essential information with a relatively small set of relevant methods, and beyond that, further context offers marginal benefits.
@@ 5.2.5 Impact of Code LLM’s Size
Figure 5: Rambo’s performance by the sizes of $\mathscr{M}$ (DeepSeek Coder)
{mdtheorem} namelabel Answer to RQ2: Rambo’s performance improves with relevant context and larger Code LLMs. Lexical retrieval provides the best trade-off between performance and efficiency. Performance gains diminish with excessive relevant context and larger models increase accuracy but slow down generation. Generating multiple sketches does not significantly enrich the set of essential code elements, instead, it may introduce redundancy, leading to reduced performance.
@@ 5.3.2 Impact of Repository Size
Fig. 7 illustrates Rambo’s performance across repositories of various sizes. Rambo performs optimally on smaller repositories (20 out of 26 repositories with fewer than 800 files). On average, it achieves a CodeBLEU score of 54.5, a Compilation Rate (CR) of 64%, and an Exact Match (EM) rate of 21.1%. However, as the repository size increases, there is a decline in performance across all metrics. In particular, for very large repositories containing more than 2,000 files, CodeBLEU drops to 49.23, and EM decreases to 13.95%. These findings suggest that while Rambo excels in smaller repositories, its ability to maintain high accuracy with larger and more complex repositories.
Figure 7: Rambo’s performance by repository sizes (files)
5.3.3 Impact of Programming Languages
Table 6: Repo-level MBC performance Ramboand the others on RepoEval dataset (Python)
BLEU CodeBLEU Pass@1 EM
Rambo Oracle 56.09 52.58 46.33 20.11
RepoCoder Oracle 52.71 48.91 40.47 15.28
Infile 35.82 34.22 24.63 5.90
RawRAG 48.21 44.73 36.66 14.21
RLCoder 52.79 50.38 43.70 18.23
RepoCoder 51.08 48.10 40.18 16.89
Rambo 55.78 52.39 46.92 20.11
Table 6 shows the performance of Rambo with Code Llama on RepoEval which is a Python benchmark proposed simultaneously with RepoCoder [15]. The results on this Python benchmark indicate that Rambo maintains significant improvements of 19%–240% in EM and up to 90% in Pass@1 over the existing approaches, even their upper-bound performance. These demonstrate Rambo’s ability to generate accurate and contextually relevant method bodies regardless of language. These findings affirm that the strategies employed by Rambo, particularly the identification of essential code elements and relevant usage extraction, are robust and effectively generalize across diverse programming environments.
{mdtheorem} namelabel Answer to RQ3: Rambo performs best with balanced surrounding context ($C_{f}$) and excels in smaller repositories. Performance declines in larger repositories due to increased complexity and generation times. Additionally, our experiments on both Java and Python demonstrate that Rambo’s approach generalizes effectively across programming languages.
@@ 5.5 Threats to Validility
Threats to internal validity include the hyperparameter settings of the baseline methods and Rambo. To mitigate this threat, we systematically varied the settings of Rambo and thoroughly examined its performance (as discussed in Section 5). Additionally, we reused the implementations and settings from the original papers of the baseline approaches [15] to ensure consistency in comparison. Another potential threat arises from the methods used to analyze generated sketch code and identify essential elements within the repository. To address this, we employed a robust code analyzer [20] and popular retrieval mechanisms applied by existing studies [15, 16, 17], ensuring our approach aligned with established practices. Furthermore, to minimize the threat of implementation errors, we conducted a detailed review of our code and have made it publicly available [26] to facilitate replication and validation by other researchers.
Threats to construct validity relate to the appropriateness of our evaluation procedure. We utilized widely accepted metrics for both text-based and code-based evaluations, including BLEU, CodeBLEU, Exact Match, Compilation Rate, and pass@k, which are standard measures in code completion and code generation research. A potential threat remains due to the absence of a human evaluation, which could provide a more nuanced assessment of our approach. To address this limitation, we plan to incorporate human evaluations in future work, wherein experienced developers will assess the generated method bodies based on criteria such as functionality, accuracy, and adherence to repository standards. Additionally, the controlled environments used in our experiments may not fully represent real-world scenarios. To mitigate this, we conducted evaluations in diverse settings and examined performance across multiple scenarios and configurations.
Threats to external validity primarily concern the generalizability of our results to different repositories and programming languages. To reduce this threat, we selected representative code LLMs models and repositories widely recognized in both NLP and SE communities [7, 19, 6]. Another potential threat is using a limited set of Java and Python repositories in our experiments. Our current study focuses on a specific set of repositories, which may limit the applicability of our findings to other types of codebases or programming languages. In future work, we intend to extend our experiments to include a broader range of repositories and languages to enhance the generalizability of our results.
@@ 6 Related Work
Rambo is closely related to the research on repository-level code completion, which utilizes the broader context provided by an entire code repository [32, 33, 34, 35, 15, 16, 18, 17]. RepoPrompts [35] introduces an approach that first generates prompts reflecting the repository’s structure and predefined relevant contexts, then learns to select the prompt most likely to produce the desired output. In the work by Wang et al. [18], the task of retrieving useful context is framed as a reinforcement learning problem, where the retriever parameters are iteratively updated. MGD [32] monitors the code generated by LLMs, queries static analysis tools in the background, and leverages the resulting information to guide the LLMs during the decoding stage. RepoCoder [15] integrates a similarity-based retriever and a code LLM into an iterative retrieval-generation pipeline. RepoHyper [17] improves the context retrieval quality over the prior work by introducing a hypergraph-based retriever that learns to model cross-file dependencies. While RepoHyper demonstrates strong performance in selecting informative context for code generation, it operates under the assumption that the necessary imported files are already present in the file being completed. Furthermore, RepoHyper requires a training phase for re-ranking and selecting suitable retrieved contexts, which may limit its applicability in dynamic or unseen repositories.
Although these efforts show promising performance, Rambo differs from them in that Rambo requires neither training nor modifications to the underlying code LLMs. To further improve its practical usability, Rambo makes no assumptions about the existence of imported files in the current file. Additionally, instead of similarity-based retrieval, Rambo emphasizes the extraction of code elements most relevant to the infilling method and their usage contexts. This design ensures that the generated method bodies are not only contextually accurate but also aligned with the repository’s specific coding practices.
@@
Recently, several learning-based approaches have been proposed for specific Software Engineering tasks such as code translation/migration [47, 48, 49, 50, 51, 52], text-code translation [53, 54, 55, 56, 57], test generation [58, 59], bug detection [60, 61, 62, 63], bug fix identification [64], and program repair [65, 66].
@@ 7 Conclusion
[…]
Acknowledgement
This research is supported by Vietnam National Foundation for Science and Technology Development (NAFOSTED) under grant number 102.03-2023.14. This research is also partly supported by OpenAI’s Researcher Access Program.